OpenStudy (anonymous):

Select the equations that are parallel and perpendicular to y = -3x - 1 and that pass through the point (3, 1).

OpenStudy (anonymous):

parallel: y = -3x + 6 perpendicular: y = 1/3x + 2 2/3 parallel: y = -3x + 10 perpendicular: y = 1/3x parallel: y = 1/3x + 1 perpendicular: y = 3x - 1 parallel: y = -1/3x perpendicular: y = -3x

OpenStudy (anonymous):

you can rule out the last two because in order to be parallel, lines MUST have the same slope. so it comes down to just the first two... and since the perpendicular lines have the correct slope, just plug in the given point to see which one gives a true statement.